Determine the radius of a shift in decimeter of a pulley rotating at 300 revolution per minute when it drives another shaft rotation at 800 revolution per minutes having a diameter of 0.2m


1
Expert's answer
2021-02-05T04:29:55-0500

The velocity "v" of the point of contact on the pitch circles is the same on both gears, and is given by


"v=r_1\omega_1=r_2\omega_2"


So,


"\frac{r_1}{r_2}=\frac{\omega_2}{\omega_1}\to\frac{r_1}{r_2}=\frac{n_2}{n_1}\to r_1=r_2\frac{n_2}{n_1}="


"=0.2\cdot\frac{800}{300}\approx0.53\ (m)" . Answer